OpenStudy (anonymous):

Can x^4 +2x-1=0 be solved algebraically?

OpenStudy (anonymous):

yes, with synthetic division

OpenStudy (anonymous):

or long division whichever you prefer

OpenStudy (anonymous):

confirm that the 2nd term is just an x if that is the case must use differnt method than factoring.

OpenStudy (anonymous):

Could I solve it by completing the square also?

OpenStudy (anonymous):

The only rational zeros possible are 1 and -1. Try them. Then use synthetic division to get a quotient with a remainder of 0

OpenStudy (anonymous):

@Dock-\[x ^{4}+2x-1=0\] This is the question, what must you divide this with?

OpenStudy (anonymous):

Well 1 or -1 don't work, it will be irrational zeros

OpenStudy (anonymous):

Awesome one Luis

OpenStudy (accessdenied):

Why do you use the quadratic formula on a quartic function?

OpenStudy (anonymous):

Got it correct! Thanks!

OpenStudy (anonymous):

you can't use quadratic formula on quartic function...

OpenStudy (anonymous):

you must have mistyped x^2 as x^4

OpenStudy (anonymous):

Sorry guys I meant x^2

OpenStudy (anonymous):

yes then use quadratic formula like luis

OpenStudy (anonymous):

\[-2\pm \sqrt{2+4}/2\]Rofl Jtla its easy then.. x^2 +2x-1=0 Use the formula: \[-2\pm \sqrt{6}/2\]